In zimbabwe the rate of inflation hit 90 sextillion percent in 2009, with prices increasing tenfold every day. At that rate, how much would a $3000 car cost two days later ?
Answer #1

The prices increase tenfold every day and the rate of inflation is 90 sixtillion

cost of a $100 car next day = 100*10 = 1000

cost of a $3000 car two days later = 3000*10^2 = 300000